OpenStudy (gabbyalicorn):

A wheel has a diameter of 21 inches. Approximately how far does it travel in 3 revolutions? A. 63 in. B. 65.94 in. C. 98.91 in. D. 197.82 in.

OpenStudy (gabbyalicorn):

@MonKiwi

OpenStudy (gabbyalicorn):

I think A

OpenStudy (gabbyalicorn):

because 21 x 3 = g3

OpenStudy (gabbyalicorn):

63* i meant

OpenStudy (anonymous):

i would think its A as well, but i would like to have it confirmed just to be sure.

OpenStudy (gabbyalicorn):

yeah, okay.

OpenStudy (gabbyalicorn):

@johnweldon1993

OpenStudy (anonymous):

A is not the answer

OpenStudy (gabbyalicorn):

Why?

OpenStudy (gabbyalicorn):

@UsukiDoll

OpenStudy (gabbyalicorn):

@unknownunknown

OpenStudy (anonymous):

Keys: C=Circumference D=Diameter R=Radius 3.14=pi How to find the circumference: D=21 Divide that by 2 = 10.5 The formula for the circumference of a circle is 2×3.14×r Plug in the numbers: 2×3.14×10.5 Which equals 65.94 Then, we have to multiply that by 3. So, 65.94 x 3 = 197.82 Which is our answer.

OpenStudy (gabbyalicorn):

@geerky42

OpenStudy (unknownunknown):

So, remember c/d = pi. Re-arrange to get c = pi * d One circumference of a circle is equivalent to a full revolution. So we multiply by 3 for 3 revolutions. pi * d * 3 = 197.82, as YoungTrilogy stated.

OpenStudy (gabbyalicorn):

okay, thank you :) √